# Gradient-flow SDEs have unique transient population dynamics

### v2

@@ Abstract
Identifying the drift and diffusion of an SDE from its population dynamics is a notoriously challenging task. Researchers in machine learning and single cell biology have only been able to prove a partial identifiability result: for potential-driven SDEs, the gradient-flow drift can be identified from temporal marginals if the Brownian diffusivity is already known. Existing methods therefore assume that the diffusivity is known a priori, despite it being unknown in practice. We dispel the need for this assumption by providing a complete characterization of identifiability: the gradient-flow drift and Brownian diffusivity are jointly identifiable from temporal marginals if and only if the process is observed outside of equilibrium. Given this fundamental result, we propose nn-APPEX, the first Schrödinger Bridge–based inference method that can simultaneously learn the drift and diffusion of gradient-flow SDEs solely from observed marginals. Extensive numerical experiments show that nn-APPEX’s ability to adjust its diffusion estimate enables accurate inference, while previous Schrödinger Bridge methods obtain biased drift estimates due to their assumed, and likely incorrect, diffusion.
@@ 1 Introduction
Mathematical biologists conceptualize the evolving genetic profiles of cells, or developmental trajectories, with Waddington’s epigenetic landscape, which likens differentiating cells to marbles rolling along a surface (Waddington, 1935). Cell development is therefore thought to be driven by the gradient of this unknown landscape. More broadly, gradient-driven dynamics describe a host of real-world processes, such as contaminant flow within groundwater systems, electric current within an electromagnetic field, and molecular dynamics driven by interatomic potential energy. Since these processes exhibit stochasticity, they are often modeled by gradient-flow SDEs, such that the drift is given by $\nabla\Psi(X_{t})$, and stochasticity is introduced by Brownian motion $W_{t}$, with diffusivity $\sigma^{2}>0$ (Weinreb et al., 2018; Lavenant et al., 2024; Lelievre and Stoltz, 2016; Yeo et al., 2021):
@@
Despite receiving significant attention from researchers in mathematical biology and machine learning (Hashimoto et al., 2016; Weinreb et al., 2018; Neklyudov et al., 2023; Lavenant et al., 2024; Yeo et al., 2021; Zhang et al., 2021; Chizat et al., 2022; Bunne et al., 2022; Shen et al., 2025; Zhang, 2024; Guan et al., 2024; Terpin et al., 2024; Persiianov et al., 2025), the most comprehensive identifiability results only prove that the drift $\nabla\Psi$ can be identified from $\bigl(p(\cdot,t)\bigr)_{t\in[0,T]}$ if the diffusivity $\sigma^{2}$ is already known:
@@
This result, popularized in the mathematical biology community by Weinreb et al. (2018) and in the machine learning community by Hashimoto et al. (2016), has informed the predominant view that the only way to achieve principled inference is to assume that the diffusion is already known. As a result of this theory, virtually all inference methods have been developed to estimate $\nabla\Psi$ given a known diffusivity $\sigma^{2}$ (Hashimoto et al., 2016; Neklyudov et al., 2023; Lavenant et al., 2024; Yeo et al., 2021; Zhang et al., 2021; Chizat et al., 2022; Bunne et al., 2022; Shen et al., 2025; Zhang, 2024). While the drift is often the principal object of interest, the diffusion is also typically unknown, and contains important insights in its own right, such as the sensitivity of cell fates to initial conditions (Forrow, 2024). Perhaps most importantly, misspecified diffusion can significantly bias drift estimation. We provide a simple example in Fig.˜1, where the true landscape is flat, but overestimated diffusion introduces a sink and underestimated diffusion introduces a source.
@@
It has remained an open question whether it is possible and, if so, under which circumstances, both the drift and diffusion of a gradient-flow SDE (1) can be identified from its marginals. Our first theoretical contribution is a complete solution to this problem. Theorem˜4.2 provides the necessary and sufficient conditions for identifying the gradient flow drift and diffusivity from marginals:
@@
We also extend this result beyond the setting of continuous observation $t\in[0,T]$, by proving in ˜4.3 that observing three distinct marginals identifies the true SDE from any countable set of candidates, with probability $1$. Since our identifiability conditions are commonly observed in practice, our results provide:
@@
With our identifiability theory guaranteeing that joint inference is well-posed, we propose nn-APPEX, the first Schrödinger Bridge (SB) based method capable of estimating both the gradient flow drift $\nabla\Psi$ and the diffusivity $\sigma^{2}$. In particular, nn-APPEX introduces a diffusion estimation step at each iteration while utilizing neural networks to flexibly infer the drift field. To assess the importance of nn-APPEX’s diffusion estimation step, we compare its performance against previous SB methods on popular benchmarks (Terpin et al., 2024; Persiianov et al., 2025). The results corroborate our theory in more practical settings, while also highlighting the importance of learning the correct diffusion for accurate drift estimation, as nn-APPEX markedly outperforms previous SB methods.
2 Related works
Our work contributes novel theory and methodology for identifying the true SDE from observed marginals. We overview relevant literature below.
Identifiability theory for population dynamics.
A popular approach for marginals-based inference is to reconstruct trajectories with entropic optimal transport, which is equivalent to solving a Schrödinger Bridge (SB) problem (Peyré et al., 2019, Prop. 4.2). The first such method was Waddington-OT (WOT) (Schiebinger et al., 2019), which performs trajectory inference by solving a single SB problem with respect to a Brownian motion reference with a prescribed diffusivity $\sigma^{2}$. WOT spawned many variations, which incorporate additional information, like cell lineages (Forrow and Schiebinger, 2021; Ventre et al., 2023), or improve robustness given limited samples (Zhang et al., 2021; Lavenant et al., 2024; Chizat et al., 2022). In the last year, SB methods were improved with the introduction of iterative reference refinement, i.e. SBIRR (Shen et al., 2025; Zhang, 2024). Rather than fixing Brownian motion as a reference, SBIRR fixes the diffusivity $\sigma^{2}$ and iteratively re-estimates the drift of the reference process. In contrast to SBIRR, the method that we introduce in this work, nn-APPEX, additionally re-estimates the diffusivity at each iteration. nn-APPEX builds upon the tri-level iterative scheme of APPEX (Guan et al., 2024), which was specifically developed for linear additive noise SDEs, by using a neural network to estimate the nonlinear drift fields encountered in gradient-flow SDEs.
Other inference methods for population dynamics.
In addition to SB methods, variational energy-based methods (Hashimoto et al., 2016; Bunne et al., 2022; Neklyudov et al., 2023; Terpin et al., 2024; Persiianov et al., 2025) have been developed for SDE inference from population dynamics. These methods leverage the fact that the marginals of potential-driven SDEs minimize a corresponding energy, and use discrete numerical schemes to approximate the underlying parameters. Of these, only JKOnet (Terpin et al., 2024), jointly estimates drift and diffusion. However, their work did not study identifiability theory, and therefore lacked guarantees for principled inference. We perform experiments comparing the performance of nn-APPEX and JKOnet in the Supplement. We also note that recent work presents another framework for joint inference via maximum mean discrepancy minimization (Berlinghieri et al., 2025).
3 Mathematical background
As is standard in stochastic analysis (Guan et al., 2024; Shen et al., 2025; Berlinghieri et al., 2025), we ensure that the gradient-flow SDE (1) is well-defined by assuming that the drift $\nabla\Psi$ satisfies Lipschitz continuity and linear growth, $\|\nabla\Psi(x)\|\leq K(1+\|x\|)$ for some $K>0$ (Oksendal, 2013, Theorem 5.5). Given these conditions, and an initial distribution $p(\cdot,0)$ with finite second moments, the population dynamics are defined by the Fokker-Planck equation. For all $x\in\mathbb{R}^{d},t\geq 0$,
@@ 3 Mathematical background
While the Fokker-Planck equation (4) may not be defined pointwise if the arguments are not sufficiently smooth, it can be defined in the weak distributional sense, such that $\mathcal{L}^{*}_{\shortminus\nabla\Psi,\sigma^{2}}$ is an operator on probability distributions (see Section˜A.1). We therefore adopt the weak formulation and use standard notation by omitting $x$ (Bogachev et al., 2022), e.g. $\mathcal{L}^{*}_{\shortminus\nabla\Psi,\sigma^{2}}\bigl(p(\cdot,t)\bigr)$ rather than $\mathcal{L}^{*}_{\shortminus\nabla\Psi,\sigma^{2}}\bigl(p(\cdot,t)\bigr)(x)$.
@@
Consider two SDEs with quadratic potential, i.e. Ornstein-Uhlenbeck processes,
@@ 4 Identifiability results
In this section, we present our main result in Theorem˜4.2, which states that the gradient-flow SDE (1) is identifiable from its marginals, $\bigl(p(\cdot,t)\bigr)_{t\in[0,T]}$, if and only if it is observed outside of equilibrium. We then show in ˜4.3 that three distinct marginals identify the true SDE with probability $1$ from any countable set of candidate SDEs.
We first prove that non-stationarity is a necessary condition for identifiability, by extending ˜3.2 for arbitrary potentials.
Proposition 4.1.
@@
$\Leftrightarrow 0$ $=\alpha\left(\nabla\cdot(p_{\mathrm{eq}}\nabla\Psi+\frac{\sigma^{2}}{2}\Delta p_{\mathrm{eq}}\right)$
@@
˜4.1 proves that if $p(\cdot,t)=p(\cdot,0)$ $\forall t\geq 0$ then the SDE is non-identifiable.
Now, we prove that non-identifiability implies that $p(\cdot,t)=p(\cdot,0)$ $\forall t\geq 0$. Suppose that SDEs with parameters $(-\nabla\Psi_{1},\sigma_{1}^{2})$ and $(-\nabla\Psi_{2},\sigma_{2}^{2})$ produce the same marginals $\bigl(p(\cdot,t)\bigr)_{t\in[0,T]}$. Without loss of generality, we may assume that $\sigma_{1}^{2}>\sigma_{2}^{2}$, because the previous partial result (2) states if $\sigma_{1}^{2}=\sigma_{2}^{2}$, then $\nabla\Psi_{1}=\nabla\Psi_{2}$ also follows (Lavenant et al., 2024, Theorem 2.1). Then, from ˜3.1, for all observed marginals $p(\cdot,t)$,
@@
We note that the PDE (10) is the Fokker-Planck equation, parametrized by the “residual” drift and diffusion. Then, to prove that $p(\cdot,t)=p(\cdot,0)$ $\forall t\geq 0$, it suffices to show that there is at most one probability distribution, $p(\cdot,t)=\mu$, which solves (10), since each marginal would then coincide. Since the residual diffusivity $\sigma_{1}^{2}-\sigma_{2}^{2}>0$ is nondegenerate, and the residual drift $\nabla(\Psi_{1}-\Psi_{2})$ is Lipschitz and obeys linear growth, it follows that the residual Fokker-Planck equation (10) has at most one stationary distribution (Bogachev et al., 2022, Theorem 4.16, Example 4.1.8). ∎
@@
While ˜4.1 shows that either the diffusivity $\sigma^{2}$ or the potential $\Psi$ needs to be known to disambiguate the true SDE from other SDEs with the same stationary distribution, Theorem˜4.2 shows that this assumption is not required if the marginals are transient. This is important for real-world inference since the condition $p(\cdot,t)=p(\cdot,0)$ $\forall t\geq 0$ is easily testable, and it is much more common to observe transient behaviour than it is to know an accurate estimate of diffusivity a priori (Forrow, 2024). However, Theorem˜4.2 still relies on continuous observation, while we only observe a finite number of marginals in practice. We address this by showing that observing three distinct marginals is enough to identify the true SDE from any countable set of candidates, e.g. the set of SDEs that can be represented on a computer. To facilitate the proof, we assume that $\Psi\in C^{\infty}(\mathbb{R}^{d})$ for this result (see details in Section˜A.2).
@@
Let $\mathcal{S}$ be a countable set of gradient-flow SDEs with smooth potentials, which all share the same initial distribution $p(\cdot,0)$. If we observe distinct marginals $\{p(\cdot,t_{i})\}_{i=1}^{3}$ from an SDE in $\mathcal{S}$, such that the measurement times $\{t_{i}\}_{i=1}^{3}$ are uniformly sampled, i.e. $t_{i}\sim\mathrm{Unif}[T_{i},T_{i+1}]$, for some $0<T_{1}<T_{2}<T_{3}<T_{4}$, then, with probability $1$, this is the only SDE in $\mathcal{S}$ with marginals $\{p(\cdot,t_{i})\}_{i=1}^{3}$.
@@
By the countability of $\mathcal{S}$, it suffices to prove that the event that any two distinct SDEs in $\mathcal{S}$ share the same marginals at each of the times $\{t_{i}\}_{i=1}^{3}$, has probability 0. Denote their parameters by $(-\nabla\Psi_{1},\sigma_{1}^{2})\neq(-\nabla\Psi_{2},\sigma_{2}^{2})$ and their marginals by $\bigl(p(\cdot,t)\bigr)_{t\geq 0}$ and $\bigl(q(\cdot,t)\bigr)_{t\geq 0}$. Suppose for contradiction that if we sample $t_{i}\sim\mathrm{Unif}[T_{i},T_{i_{+}1}]$, then we observe $p(\cdot,t_{i})=q(\cdot,t_{i})$ $\forall i\in\{1,2,3\}$, with some nonzero probability.
Since we sample $t_{i}\sim\mathrm{Unif}[T_{i},T_{i_{+}1}]$, then the set of coincidence times in each subinterval, $\mathcal{I}_{i}=\{t\in[T_{i},T_{i_{+}1}]\mid p(\cdot,t)=q(\cdot,t)\}$, must be infinite to ensure nonzero probability. By Bolzano-Weierstrass, we may construct the convergent sequence $t_{1}^{(n)}\to t_{1}^{*}\in[T_{1},T_{2}]$, using times $t_{1}^{(n)}\in\mathcal{I}_{1}$, and the convergent sequence $t_{2}^{(n)}\to t_{2}^{*}\in[T_{2},T_{3}]$, using times $t_{2}^{(n)}\in\mathcal{I}_{2}$. At the limit points, $t_{1}^{*}$ and $t_{2}^{*}$, we have $p(\cdot,t_{i}^{*})=q(\cdot,t_{i}^{*})$ and $\frac{\partial}{\partial_{t}}p(\cdot,t_{i}^{*})=\frac{\partial}{\partial_{t}}q(\cdot,t_{i}^{*})$ (see Lemma˜A.2 and the proof of Hashimoto et al. (2016, Corollary 2)). Hence,
@@
It then follows from Theorem˜4.2 that (11) and (12) are both solved by at most one distribution $\mu$. Thus, $\mu=p(\cdot,t_{1}^{*})=p(\cdot,t_{2}^{*})$. However, by ˜A.3, the only marginal that can repeat at distinct times is the stationary Gibbs distribution $p_{\mathrm{eq}}$. This follows from a variation of Boltzmann’s H-theorem, which states that the marginals of a gradient-flow SDE (1) decrease free energy, which is uniquely minimized at $p_{\mathrm{eq}}$ (Jordan et al., 1998; Chafaï, 2015). Both processes must therefore reach $p_{\mathrm{eq}}$ by $t_{1}^{*}$, such that $p(\cdot,t)=q(\cdot,t)=p_{\mathrm{eq}}$ for all $t\geq t_{1}^{*}$. In particular, since $t_{1}^{*}$ is in the first subinterval, we have $p(\cdot,t_{2})=p(\cdot,t_{3})=p_{\mathrm{eq}}$, which contradicts the assumption that the marginals $\{p(\cdot,t_{i})\}_{i=1}^{3}$ are distinct. ∎

### v3

@@ Abstract
Identifying the drift and diffusion of an SDE from its population dynamics is a notoriously challenging task. Researchers in machine learning and single-cell biology have only been able to prove a partial identifiability result: for potential-driven SDEs, the gradient-flow drift can be identified from temporal marginals if the Brownian diffusivity is already known. Existing methods therefore assume that the diffusivity is known a priori, despite it being unknown in practice. We dispel the need for this assumption by providing a complete characterization of identifiability: the gradient-flow drift and Brownian diffusivity are jointly identifiable from temporal marginals if and only if the process is observed outside of equilibrium. Given this fundamental result, we propose nn-APPEX, the first Schrödinger Bridge–based inference method that can simultaneously learn the drift and diffusion of a gradient-flow SDE solely from observed marginals. Extensive experiments show that nn-APPEX’s ability to adjust its diffusion estimate enables accurate inference, while previous Schrödinger Bridge methods obtain biased drift estimates due to their assumed, and likely incorrect, diffusion.
@@ 1 INTRODUCTION
Mathematical biologists conceptualize the evolving genetic profiles of cells, or developmental trajectories, with Waddington’s epigenetic landscape, which likens differentiating cells to marbles rolling along a surface (Waddington, 1935). Cell development is therefore thought to be driven by the gradient of this unknown landscape. More broadly, gradient-driven dynamics describe a host of real-world processes, such as contaminant flow within groundwater systems, electric current within an electromagnetic field, and molecular dynamics driven by interatomic potential energy. Since these processes exhibit stochasticity, they are often modeled by gradient-flow stochastic differential equations (SDEs), such that the drift is given by $\nabla\Psi(X_{t})$, and stochasticity is introduced by Brownian motion $W_{t}$, with diffusivity $\sigma^{2}>0$ (Weinreb et al., 2018; Lavenant et al., 2024; Lelievre and Stoltz, 2016):
@@
Despite receiving significant attention from researchers in mathematical biology and machine learning (Hashimoto et al., 2016; Weinreb et al., 2018; Neklyudov et al., 2023; Lavenant et al., 2024; Yeo et al., 2021; Zhang et al., 2021; Chizat et al., 2022; Bunne et al., 2022; Shen et al., 2025; Terpin et al., 2024), the most comprehensive identifiability results only prove that the drift $\nabla\Psi$ can be identified from $\bigl(p(\cdot,t)\bigr)_{t\in[0,T]}$ if the diffusivity $\sigma^{2}$ is already known:
@@
This result, popularized in the mathematical biology community by Weinreb et al. (2018) and in the machine learning community by Hashimoto et al. (2016), has informed the predominant view that the only way to achieve principled inference is to assume that the diffusion is already known. As a result of this theory, virtually all inference methods have been developed to estimate $\nabla\Psi$ given a known diffusivity $\sigma^{2}$ (Hashimoto et al., 2016; Neklyudov et al., 2023; Lavenant et al., 2024; Yeo et al., 2021; Zhang et al., 2021; Chizat et al., 2022; Bunne et al., 2022; Shen et al., 2025; Zhang, 2024). While the drift is often the principal object of interest, the diffusion is also typically unknown, and contains important insights in its own right, such as the sensitivity of cell fates to initial conditions (Forrow, 2024). Perhaps most importantly, misspecified diffusion can significantly bias drift estimation. We provide a simple example in Figure˜1, where the true landscape is flat, but overestimated diffusion introduces a sink and underestimated diffusion introduces a source.
@@
It has remained an open question whether it is possible and, if so, under which circumstances, both the drift and diffusion of a gradient-flow SDE (1) can be identified from its marginals. Our first theoretical contribution is a complete solution to this problem. Theorem˜3.2 provides the necessary and sufficient conditions for identifying the gradient flow drift and diffusivity from marginals:
@@
We also extend this result beyond the setting of continuous observation $t\in[0,T]$, by proving in ˜3.3 that observing three distinct marginals identifies the true SDE from any countable set of candidates, with probability $1$. Since our identifiability conditions are commonly observed in practice, our results provide:
@@
With our identifiability theory guaranteeing that joint inference is well-posed, we propose nn-APPEX, the first Schrödinger Bridge (SB) based method capable of estimating both the gradient flow drift $\nabla\Psi$ and the diffusivity $\sigma^{2}$. In particular, nn-APPEX introduces a diffusion estimation step at each iteration while utilizing a neural network to flexibly infer the drift field. To assess the importance of nn-APPEX’s diffusion estimation step, we compare its performance against previous SB methods on simulated benchmark data (Terpin et al., 2024; Persiianov et al., 2025) and on a single-cell dataset from human embryonic stem cells (Chu et al., 2016; Shen et al., 2025). The results corroborate our identifiability theory, while also highlighting the importance of learning the correct diffusion for accurate drift estimation, as nn-APPEX markedly outperforms previous SB methods.
2 MATHEMATICAL SETUP
As is standard in stochastic analysis (Shen et al., 2025; Berlinghieri et al., 2025), we ensure that the gradient-flow SDE (1) is well-defined by assuming that the drift $\nabla\Psi$ satisfies Lipschitz continuity and linear growth, $\|\nabla\Psi(x)\|\leq K(1+\|x\|)$ for some $K>0$ (Oksendal, 2013, Theorem 5.5). Given these conditions, and an initial distribution $p(\cdot,0)$ with finite second moments, the population dynamics are defined by the Fokker-Planck equation. For all $x\in\mathbb{R}^{d},t\geq 0$,
@@ 2 MATHEMATICAL SETUP
While the Fokker-Planck equation (4) may not be defined pointwise if the arguments are not sufficiently smooth, it can be defined in the weak distributional sense, such that $\mathcal{L}^{*}_{\shortminus\nabla\Psi,\sigma^{2}}$ is an operator on probability distributions (see Section˜A.1). We therefore adopt the weak formulation and use standard notation by omitting $x$ (Bogachev et al., 2022), e.g., $\mathcal{L}^{*}_{\shortminus\nabla\Psi,\sigma^{2}}\bigl(p(\cdot,t)\bigr)$ rather than $\mathcal{L}^{*}_{\shortminus\nabla\Psi,\sigma^{2}}\bigl(p(\cdot,t)\bigr)(x)$.
@@
Consider two SDEs with quadratic potential, i.e., Ornstein-Uhlenbeck processes,
@@ 3 IDENTIFIABILITY RESULTS
In this section, we present our main result in Theorem˜3.2, which states that the gradient-flow SDE (1) is identifiable from its marginals, $\bigl(p(\cdot,t)\bigr)_{t\in[0,T]}$, if and only if it is observed outside of equilibrium. We then show in ˜3.3 that three distinct marginals identify the true SDE with probability $1$ from any countable set of candidate SDEs.
We first prove that non-stationarity is a necessary condition for identifiability, by extending ˜2.2 for arbitrary potentials.
Proposition 3.1 (Non-identifiability from equilibrium).
@@
$\Leftrightarrow 0$ $=\alpha\left(\nabla\cdot(p_{\mathrm{eq}}\nabla\Psi+\frac{\sigma^{2}}{2}\Delta p_{\mathrm{eq}})\right)$
@@
˜3.1 proves that if $p(\cdot,t)=p(\cdot,0)$ $\forall t\geq 0$ then the SDE is non-identifiable.
Now, we prove that non-identifiability implies that $p(\cdot,t)=p(\cdot,0)$ $\forall t\geq 0$. Suppose that SDEs with parameters $(-\nabla\Psi_{1},\sigma_{1}^{2})$ and $(-\nabla\Psi_{2},\sigma_{2}^{2})$ produce the same marginals $\bigl(p(\cdot,t)\bigr)_{t\in[0,T]}$. Without loss of generality, we can assume that $\sigma_{1}^{2}>\sigma_{2}^{2}$, because the previous partial result (2) states that if $\sigma_{1}^{2}=\sigma_{2}^{2}$, then $\nabla\Psi_{1}=\nabla\Psi_{2}$ also follows (Lavenant et al., 2024, Theorem 2.1). Then, from ˜2.1, for all observed marginals $p(\cdot,t)$,
@@
We note that the PDE (10) is the Fokker-Planck equation, parametrized by the “residual” drift and diffusion. Then, to prove that $p(\cdot,t)=p(\cdot,0)$ $\forall t\geq 0$, it suffices to show that there is at most one probability distribution, $p(\cdot,t)=\mu$, which solves (10), since each marginal would then coincide. Since the residual diffusivity $\sigma_{1}^{2}-\sigma_{2}^{2}>0$ is nondegenerate, and the residual drift $\nabla(\Psi_{1}-\Psi_{2})$ is Lipschitz and obeys linear growth, it follows that the residual Fokker-Planck equation (10) has at most one stationary distribution (Bogachev et al., 2022, Theorem 4.1.6, Example 4.1.8). ∎
@@
While ˜3.1 shows that either the diffusivity $\sigma^{2}$ or the potential $\Psi$ needs to be known to disambiguate the true SDE from other SDEs with the same stationary distribution, Theorem˜3.2 shows that this assumption is not required if the marginals are transient. This is important for real-world inference since the condition $p(\cdot,t)=p(\cdot,0)$ $\forall t\geq 0$ is easily verifiable, and it is much more common to observe transient behaviour than it is to know an accurate a priori estimate of the diffusivity coefficient (Forrow, 2024). However, Theorem˜3.2 still relies on continuous observation, while we only observe a finite number of marginals in practice. We address this by showing that observing three distinct marginals is enough to identify the true SDE from any countable set of candidates, e.g. the set of SDEs that can be represented on a computer. To facilitate the proof, we assume that $\Psi\in C^{\infty}(\mathbb{R}^{d})$ (see details in Section˜A.2).
@@
Let $\mathcal{S}$ be a countable set of gradient-flow SDEs with smooth potentials, which all share the same initial distribution $p(\cdot,0)$. If we observe distinct marginals $\{p(\cdot,t_{i})\}_{i=1}^{3}$ from an SDE in $\mathcal{S}$, such that the measurement times $\{t_{i}\}_{i=1}^{3}$ are uniformly sampled, i.e., $t_{i}\sim\mathrm{Unif}[T_{i},T_{i+1}]$, for some $0<T_{1}<T_{2}<T_{3}<T_{4}$, then, with probability $1$, this is the only SDE in $\mathcal{S}$ with marginals $\{p(\cdot,t_{i})\}_{i=1}^{3}$.
@@
By the countability of $\mathcal{S}$, it suffices to prove that the event that any two distinct SDEs in $\mathcal{S}$ share the same marginals at each of the times $\{t_{i}\}_{i=1}^{3}$ has probability 0. Denote their parameters by $(-\nabla\Psi_{1},\sigma_{1}^{2})\neq(-\nabla\Psi_{2},\sigma_{2}^{2})$ and their marginals by $\bigl(p(\cdot,t)\bigr)_{t\geq 0}$ and $\bigl(q(\cdot,t)\bigr)_{t\geq 0}$. Suppose for contradiction that if we sample $t_{i}\sim\mathrm{Unif}[T_{i},T_{i+1}]$, then we observe $p(\cdot,t_{i})=q(\cdot,t_{i})$ $\forall i\in\{1,2,3\}$, with some nonzero probability.
Since we sample $t_{i}\sim\mathrm{Unif}[T_{i},T_{i+1}]$, then the set of coincidence times in each subinterval, $\mathcal{I}_{i}=\{t\in[T_{i},T_{i+1}]\mid p(\cdot,t)=q(\cdot,t)\}$, must be infinite to ensure nonzero probability. By Bolzano-Weierstrass, we can construct the convergent sequence $t_{1}^{(n)}\to t_{1}^{*}\in[T_{1},T_{2}]$, using times $t_{1}^{(n)}\in\mathcal{I}_{1}$, and the convergent sequence $t_{2}^{(n)}\to t_{2}^{*}\in[T_{2},T_{3}]$, using times $t_{2}^{(n)}\in\mathcal{I}_{2}$. At the limit points, $t_{1}^{*}$ and $t_{2}^{*}$, we have $p(\cdot,t_{i}^{*})=q(\cdot,t_{i}^{*})$ and $\frac{\partial}{\partial_{t}}p(\cdot,t_{i}^{*})=\frac{\partial}{\partial_{t}}q(\cdot,t_{i}^{*})$ (see Lemma˜A.2 and the proof of Hashimoto et al. (2016, Corollary 2)). Hence,
@@
It then follows from Theorem˜3.2 that (11) and (12) are both solved by at most one distribution $\mu$. Thus, $\mu=p(\cdot,t_{1}^{*})=p(\cdot,t_{2}^{*})$. However, by ˜A.3, the only marginal that can repeat at distinct times is the stationary Gibbs distribution $p_{\mathrm{eq}}$. This follows from a variation of Boltzmann’s H-theorem, which states that the marginals of a gradient-flow SDE (1) decrease free energy, which is uniquely minimized at $p_{\mathrm{eq}}$ (Jordan et al., 1998; Chafaï, 2015). Both processes must therefore reach $p_{\mathrm{eq}}$ by $t_{1}^{*}$, such that $p(\cdot,t)=q(\cdot,t)=p_{\mathrm{eq}}$ for all $t\geq t_{1}^{*}$. In particular, since $t_{1}^{*}$ is in the first subinterval, we have $p(\cdot,t_{2})=p(\cdot,t_{3})=p_{\mathrm{eq}}$, which contradicts the assumption that the marginals $\{p(\cdot,t_{i})\}_{i=1}^{3}$ are distinct. ∎
4 NN-APPEX: THREE-STAGE SB REFINEMENT
We have shown in Section˜3 that the full gradient-flow SDE is identifiable from $N\geq 3$ distinct marginals, $\bigl(p(\cdot,t_{i})\bigr)_{i=0}^{N-1}$. Building on this, we propose Neural Network-based Alternating Projection Parameter Estimation from $X_{0}$ (nn-APPEX). APPEX searches for the true parameters $(-\nabla\Psi,\sigma^{2})$ by alternating between trajectory inference, drift estimation, and diffusion estimation. nn-APPEX improves upon APPEX (Guan et al., 2024) by replacing the linearly parameterized drift estimation step with a neural network, such that it can infer SDEs with nonlinear drift, e.g. arbitrary gradient-flow $\nabla\Psi$.
The three-stage optimization proceeds as follows. First, given a reference SDE $Q$, we solve a multi-marginal Schrödinger Bridge problem to infer a law on paths $P$. In particular, $P$ minimizes $\mathrm{KL}(P\|Q)$, while obeying the multi-marginal constraints, i.e., $P\in\Pi\bigl(p(\cdot,t_{i})_{i=0}^{N-1}\bigr)$. Second, from these inferred trajectories, we perform a maximum likelihood estimation (MLE) of the drift $\nabla\Psi$. For generality and tractability, we optimize over neural network parameters $\theta$. Third, from the inferred trajectories and the estimated drift, we estimate the diffusivity $\sigma^{2}$ with a closed form MLE expression. We then use the new drift and diffusion estimates to update the reference SDE $Q$ for the next iteration. Formally, at iteration $k$, nn-APPEX performs the steps
$P_{k}$ $=\mathop{\arg\min}_{P\in\Pi\left(p(\cdot,t_{i})_{i=0}^{N-1}\right)}\mathrm{KL}(P\|Q_{\shortminus\nabla\Psi_{\theta_{k-1}},\sigma^{2}_{k-1}}),$ (13)
$\theta_{k}$ $=\mathop{\arg\min}_{\theta}\mathrm{KL}(P_{k}\|Q_{\shortminus\nabla\Psi_{\theta},\sigma^{2}_{k-1}}),$ (14)
$\sigma^{2}_{k}$ $=\mathop{\arg\min}_{\sigma^{2}}\mathrm{KL}(P_{k}\|Q_{\shortminus\nabla\Psi_{\theta_{k}},\sigma^{2}}),$ (15)
and we iterate until convergence, or until a maximum number of iterations is reached. We emphasize that nn-APPEX does not require prior knowledge, since any gradient-flow SDE (1) can be used as the initial reference $Q_{-\nabla\Psi_{\theta_{0}},\sigma^{2}_{0}}$. While it is known that the $\mathrm{KL}$ divergence on the space $C([0,T],\mathbb{R}^{d})$ is infinite between SDEs with different diffusions (Shen et al., 2025), we note that given finite observation times, $\{t_{i}\}_{i=0}^{N-1}$, we can instead evaluate the $\mathrm{KL}$ divergence on $C(\{t_{i}\}_{i=0}^{N-1},\mathbb{R}^{d})$, computed over the observed couplings $p_{t_{i},t_{i+1}}$. As shown in Guan et al. (2024, Section 4) and Appendix˜D, this objective is finite for gradient-flow SDEs (1), and thus enables principled three-stage optimization. We present pseudocode in Algorithm˜1.